Producer production function is Y = 8X+2X²-4/3X³ Price of the product is 5 per unit and price of the input is 25 per unit. What is the level of output at the equilibrium level of production?


1
Expert's answer
2020-12-24T13:15:04-0500
"y'=8+4x-4x^2"


"-4x^2+4x+8=0"


"x=2"


"y(2)=16+8-\frac{32}{3}=13\frac{1}{3}"